A zero coupon bond issued by Corporation A has a residual maturity of 2 years, an expected return in continuous compounding equal to 3.12% and a face value of 1,000 USD. Knowing that the corporate bond is risky (hence the company can default) but bears no systematic risk, what is the risk-free rate in discrete annual compounding for a two-year investment?

(a) 3.17%
(b) 3.12%
(c) 3.00%
(d) Since the bond is risky it cannot be determined.

The correct answer is A
(Thumb up for correct step-by-step solution. Many thanks)
Answer #1
Price with continuous compounding = 1000/e^(0.0312*2) = 1000/1.06439 = $             939.51
Risk free rate with discrete compounding = (1000/939.51)^(1/2)-1 = 3.17%
Answer: Option [A]